# A Quantile Nelson-Siegel model

### v1

@@ 4 Empirical Application
This section is devoted to studying the effectiveness of the proposed TVP-QR-NS model for zero-coupon US treasury yields. In particular, we split the following section into an in-sample analysis of the yield curve factor and a forecasting exercise.
@@ 4.1 Data
Our empirical investigation employs zero-coupon US Treasury yields as constructed by Liu and Wu (2021) through a non-parametric kernel smoothing methodology. The application of this novel yield curve dataset by Liu and Wu (2021) is demonstrated to offer a more precise depiction of the underlying data when compared to the alternative metric presented by Gürkaynak et al. (2007). Following the methodology established by Diebold et al. (2006), our analysis encompasses zero-coupon US Treasury maturities spanning 3, 6, 9, 12, 15, 18, 21, 24, 30, 36, 48, 60, 72, 84, 96, 108, and 120 months. Additionally, our investigation includes key monthly macroeconomic indicators — Industrial Production, CPI Inflation, and the Fed Funds Rate — in line with the framework suggested by Diebold et al. (2006). These macroeconomic variables are chosen to capture the dynamic aspects of the macroeconomy and are all sourced from the US FRED database. Both Industrial Production and CPI Inflation were transformed to log-differenced growth rates. Our study period extends from November 1971 to December 2022.
@@ 4.2 In-Sample Results
Previous studies, including Diebold and Li (2006); Diebold et al. (2006, 2008); Mumtaz and Surico (2009); Mönch (2012), focused on the yield curve and its effects within a conditional mean framework. Our study is distinct from the former as it is the first to explore the yield curve from a quantile or distributional perspective. This section presents our empirical in-sample results. The first subsection details the posterior estimates of yield curve factors for each quantile level. The second subsection analyses the distributional features of short- and long-term interest rates during recent crises (the Great Recession and the COVID-19 pandemic). Lastly, the third subsection investigates the interaction between yield curve factors and the macroeconomy from a quantile perspective.
@@ 4.2.1 Quantile Yield Curve Factor Estimates
Figures 2, 3, and 4 depict the posterior mean estimates of time-varying yield curve factors at five quantiles (10%, 25%, 50%, 75%, 90%), providing insights into their dynamic evolution. A consistent finding across all graphs is the presence of substantial temporal variation. Beginning with the level factor, interpreted as the long-term interest rate (i.e., yield as maturity approaches infinity), a steady decline in this rate is discernible across all quantiles since the oil price crisis of the 1980s. Figure 2 reveals a distinct pattern: following both the dot-com recession in the early 2000s and the Great Recession of 2008-09, the posterior estimates for the 90th and median quantiles closely align, while those for the 10th quantile diverge. Conversely, after the COVID-19 recession, the opposite trend emerged, with the 10th and median quantiles exhibiting proximity while the 90th quantile diverged. This evidence suggests that post the early 2000s recession and the Great Recession, the distribution of the long-term interest rate skewed leftwards. In contrast, after the COVID-19 recession, the distribution skewed rightwards, likely attributed to the pronounced inflationary pressures experienced in the aftermath of the COVID-19 period.
@@
Turning to the slope and curvature of the yield curve in Figures 3 and 4, the degree of uncertainty associated with these factors varies over time. Notably, the curvature factor exhibits higher volatility across all quantiles than its slope counterpart. [^1] Notice that crossing of the estimated conditional quantiles for the curvature factor is not harmful since it does not have a direct link to one particular yield. A distinct pattern is evident before 1985, where the posterior estimates for the slope factor across all five quantiles are closely clustered, indicating a period of low uncertainty. In contrast, during the oil price crisis of the 1980s, the curvature factor displayed substantial dispersion across quantiles.
@@
4.2.2 The Distributional Features of the Short- and Long-Term Interest Rate
The proposed novel framework provides a distinct advantage by facilitating the investigation of distributional characteristics in both short- and long-term interest rates over time. As mentioned earlier, the long-term interest rate is identified as the level factor, while the short-term interest rate is a composite of the level and slope factors. We adhere to the methodology outlined by Mitchell et al. (2022) to construct the distributions of these interest rates. Employing an equal weighting scheme, we aggregate all MCMC draws across five quantiles and apply the Epanechnikov kernel, as proposed by Gaglianone and Lima (2012).
@@ 4.2.2 The Distributional Features of the Short- and Long-Term Interest Rate
Focusing on the short-term interest rate first, discernible evidence of a pronounced distributional shift is observed across the three selected months for both recessions. Notably, during the Great Recession (September 2009), a manifestation of multi-modality in the distribution is apparent. Additionally, after the COVID-19 recession (December 2020), the distribution became less symmetric and more negatively skewed to the left. Turning to the long-term interest rate, the evidence of distribution shifts is more prominent during the COVID-19 recession than during the Great Recession. The long-term interest rate distribution exhibits increased leftward skewness during the Great Recession (September 2009) and the subsequent period (December 2009). Conversely, during the COVID-19 recession, the distribution becomes more positively skewed to the right. Therefore, our results highlight that both short- and long-term interest rate distributions display heightened skewness and asymmetric shapes during recessionary periods. Thus highlighting the importance of using quantile models instead of standard Nelson-Siegel models that are not able to detect the left and right tails of the distribution.
4.2.3 The Interaction between the Yield Curve and Macroeconomy - Quantile VAR Parameter estimates
The proposed framework, as delineated in eq. (9), introduces an avenue for interaction between yield curve factors and macroeconomic variables by incorporating a VAR-like state equation. The estimated off-diagonal VAR coefficients for various quantile levels are presented in Figure 6, shedding light on the directional effects between yield curve factors and macroeconomic variables.
Panel A
Panel B
Figure 6: Posterior mean and credible intervals for the VAR coefficients across different quantile levels.
Panel A of the figure conveys the directional impact of yield curve factors on macroeconomic variables, while Panel B reports the converse scenario. The patterns of influence of each yield factor on the three macroeconomic variables exhibit remarkable similarity. Precisely, evidence of a U-shaped effect across quantiles is discerned for industrial production and inflation, with an inverse relationship observed for the Fed funds rate. Conversely, the impact of macroeconomic variables on the yield curve factor demonstrates higher heterogeneity. Industrial production exerts a significant positive effect on both level and curvature, whereas its impact on the slope is mixed. The effects of inflation exhibit qualitative similarity, albeit with a smaller magnitude on the level and a higher impact on the slope. In contrast, the Fed funds rate positively influences all yield curve factors. Analysing the relationship with quantile levels, the effects on the slope exhibit a U-shaped pattern for industrial production and inflation, while the Fed funds rate demonstrates an inverse U-shaped relationship. Meanwhile, the relationship with the level remains constant across quantiles. In summary, our findings provide substantial evidence that the interaction between yield curve factors and the macroeconomy significantly varies across quantiles.
4.3 Forecasting Yields Tail Risk
We conduct a pseudo-out-of-sample forecasting exercise to assess the utility of our novel quantile Nelson-Siegel approach, comparing it against the corresponding standard Nelson-Siegel conditional mean counterpart for forecasting tail risks in zero-coupon yields. The initial holdout period spans from November 1971 to January 2000, with a subsequent forecast evaluation period extending from February 2000 to December 2022 (comprising 264 evaluation periods). We generate 1st, 6th, and 12th step-ahead tail risk forecasts and employ the quantile score (QS) (see Giacomini and Komunjer, 2005; Carriero et al., 2022) as the forecast evaluation metric to gauge the performance of the two models. Following Gneiting and Ranjan (2011), we define the QS as:
Here, $QS_{\tau,i,t}$ denotes the predictive quantile of the $m$-th yield, and $\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}$ takes the value 1 if the realised value is at or below the predictive quantile and 0 otherwise. We evaluate the QS in the upper and lower tails by setting $\tau=0.9$ and $\tau=0.1$, respectively.
Table 1 presents the 10th and 90th percentiles of QS for the out-of-sample forecasting exercise. The top panel of Table 1 displays the average QS of our proposed approach relative to their corresponding conditional mean counterparts for each of the 17 yields across the evaluation period. A quantile score less than 1 implies that our proposed approach outperforms the corresponding conditional mean counterpart. Notably, our proposed approach consistently produces superior tail forecasts for all 17 yields and across all forecast horizons compared to their corresponding conditional mean counterparts. We further assess the significance of our forecasting results through a Diebold-Mariano test, with results presented in the bottom panel of Table 1, clearly indicating the statistical significance of our forecasting results in favour of our proposed approach.
Additionally, we depict the average raw QS across the 17 yields for both models over time in Figure 7. The top and bottom panels show the 10th and 90th percentiles of QS, respectively. Three key features emerge from this figure. First, reinforcing the results in Table 1, our proposed quantile Nelson-Siegel model consistently outperforms their corresponding conditional mean counterparts across all horizons and quantile levels, with the evidence being stronger for 1-step ahead forecasts. Second, the 1-step ahead QS (for both 10th and 90th percentiles) for the conditional mean model exhibits significantly more instability over time than our quantile approach. Finally, we find strong evidence in favour of the proposed time-varying parameters quantile Nelson-Siegel model for forecasting the left tail of the yield distribution, suggesting its ability to more appropriately capture the lower bound of the yield curve compared to standard conventional models.
Figure 7: Quantile Score over time for the 1-, 6-, and 12-steps ahead forecast horizons, at quantile levels 10% (top plots) and 90% (bottom plots), for the QR-NS model (blue) and the standard NS model (red). Shaded areas represent NBER recession periods.
Table 1: Forecasting results for the quantile NS relative to standard NS model, together with Diebold-Mariano test. Out-of-sample evaluation period: 2000 - 2022.
Ratio of quantile scores (quantile NS over standard NS)
10% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 0.28 0.34 0.39 0.43 0.44 0.45 0.45 0.45 0.43 0.42 0.41 0.42 0.45 0.50 0.54 0.60 0.65 0.45
6-step ahead 0.61 0.63 0.64 0.65 0.65 0.65 0.64 0.64 0.63 0.61 0.59 0.58 0.57 0.57 0.57 0.58 0.59 0.61
12-step ahead 0.67 0.68 0.69 0.70 0.70 0.71 0.71 0.70 0.70 0.69 0.67 0.66 0.65 0.63 0.63 0.62 0.62 0.67
90% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 0.47 0.55 0.56 0.55 0.53 0.51 0.50 0.49 0.49 0.49 0.50 0.53 0.54 0.56 0.57 0.57 0.56 0.53
6-step ahead 0.98 0.98 0.96 0.93 0.90 0.88 0.86 0.84 0.80 0.77 0.73 0.71 0.69 0.69 0.69 0.69 0.68 0.81
12-step ahead 0.96 0.96 0.95 0.93 0.92 0.91 0.89 0.88 0.86 0.84 0.82 0.80 0.79 0.78 0.77 0.76 0.75 0.86
Diebold-Mariano test, significance level: 1 (10%), 2 (5%), 3 (1%)
10% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
6-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
12-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
90% QS
Horizon 3 m 6 m 9 m 12 m 15 m 18 m 21 m 24 m 30 m 36 m 48 m 60 m 72 m 84 m 96 m 108 m 120 m Average
1-step ahead 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00
6-step ahead 0.00 0.00 2.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 2.59
12-step ahead 1.00 1.00 2.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 3.00 2.71

### v2

@@ 4 Empirical Application
This section focuses on evaluating the performance of the proposed TVP-QR-NS model applied to zero-coupon US Treasury yields. Specifically, the analysis is divided into two parts: an out-of-sample forecasting exercise and an in-sample assessment.
@@ 4.1 Data
Our empirical investigation employs zero-coupon US Treasury yields as constructed by Liu and Wu (2021) through a non-parametric kernel smoothing methodology. The application of this novel yield curve dataset by Liu and Wu (2021) is demonstrated to offer a more precise depiction of the underlying data when compared to the alternative metric presented by Gürkaynak et al. (2007). Following the methodology established by Diebold et al. (2006), our analysis encompasses zero-coupon US Treasury maturities spanning 3, 6, 9, 12, 15, 18, 21, 24, 30, 36, 48, 60, 72, 84, 96, 108, and 120 months. Additionally, our investigation includes key monthly macroeconomic indicators — Industrial Production, CPI Inflation, and the Fed Funds Rate — in line with the framework suggested by Diebold et al. (2006). These macroeconomic variables are chosen to capture the dynamic aspects of the macroeconomy and are all sourced from the US FRED database. Both Industrial Production and CPI Inflation were transformed to log-differenced growth rates. Our study period extends from November 1971 to December 2024.
4.2 Forecasting Yields Tail Risk
To validate the utility of our proposed framework, we conduct a pseudo-out-of-sample forecasting exercise to evaluate the performance of the quantile Nelson-Siegel approach in capturing tail risks in zero-coupon yields. The benchmark model employed for comparison is a univariate autoregressive model with one lag. Our proposed model’s performance is assessed relative to this baseline as well as three alternative conditional mean models, as detailed in Table 1. These competing specifications include: (i) a time-varying parameter model with stochastic volatility (SV); (ii) the large Bayesian vector autoregression (VAR) with common stochastic volatility (CSV) developed by Chan (2020); and (iii) the conditional mean formulation of the dynamic Nelson-Siegel model of Diebold et al. (2006).
Table 1: Overview of Competing Models
Model Description
AR(1) Autoregressive model with one lag (Benchmark model).
TVP-AR(1)-SV Time-varying parameter autoregressive model with SV and one lag.
BVAR(1)-CSV Large Bayesian VAR with one lag and common stochastic volatility.
DNS Dynamic Nelson-Siegel model of Diebold et al. (2006).
TVP-QNS Proposed Quantile-based time-varying Nelson-Siegel model.
The initial holdout period spans from November 1971 to January 2000, with a subsequent forecast evaluation period extending from February 2000 to December 2024. We generate 1st, 12th, and 36th month-ahead tail risk forecasts and evaluate the predictive performance of the competing models using the quantile score (QS) (see Giacomini and Komunjer, 2005; Carriero et al., 2022). Following Gneiting and Ranjan (2011), we define the QS as:
Here, $\mathcal{Q}_{\tau,i,t}$ denotes the predictive quantile of the $m$-th yield, and $\mathbb{I}\{y_{t}(m)\leq\mathcal{Q}_{\tau,i,t}\}$ takes the value 1 if the realised value is at or below the predictive quantile and 0 otherwise. We evaluate the QS in the upper and lower tails by setting $\tau=0.9$ and $\tau=0.1$, respectively.
Table 2: Average 10% (Lower-Tail) Quantile Scores for Selected Maturities Relative to an AR(1) Benchmark
Model 3m $p_{\text{MCS}}$ 6m $p_{\text{MCS}}$ 12m $p_{\text{MCS}}$ 24m $p_{\text{MCS}}$ 36m $p_{\text{MCS}}$ 60m $p_{\text{MCS}}$ 120m $p_{\text{MCS}}$ Avg.
1 Month-Ahead Forecast
TVP-AR(1)-SV 0.43 0.00 0.42 0.01 0.50 0.03 0.63 0.02 0.72 0.00 0.81 0.00 0.90 0.00 0.68
BVAR(1)-CSV 0.28 1.00** 0.31 1.00** 0.38 1.00** 0.50 1.00** 0.57 0.64** 0.63 0.23* 0.65 1.00** 0.52
DNS 1.72 0.00 1.32 0.00 1.06 0.00 1.13 0.00 1.23 0.00 1.29 0.00 1.07 0.00 1.20
TVP-QNS 0.48 0.00 0.46 0.00 0.47 0.04 0.53 0.46** 0.55 1.00** 0.57 1.00** 0.70 0.41** 0.55
12 Month-Ahead Forecast
TVP-AR(1)-SV 1.60 0.00 1.27 0.00 1.39 0.00 1.27 0.00 1.07 0.00 0.94 0.00 0.88 0.00 1.15
BVAR(1)-CSV 0.73 0.18* 0.77 0.18* 0.82 0.17* 0.90 0.18* 0.93 0.18* 0.94 0.18* 0.91 0.18* 0.88
DNS 1.15 0.00 1.10 0.00 1.05 0.00 1.08 0.00 1.11 0.00 1.13 0.00 1.04 0.00 1.09
TVP-QNS 0.76 1.00** 0.75 1.00** 0.73 1.00** 0.76 1.00** 0.76 1.00** 0.74 1.00** 0.64 1.00** 0.73
36 Month-Ahead Forecast
TVP-AR(1)-SV 0.00 0.00 0.00 0.00 0.00 0.00 2.84 0.00
BVAR(1)-CSV 1.13 0.00 1.17 0.00 1.19 0.00 1.24 0.00 1.26 0.00 1.27 0.00 1.25 0.00 1.23
DNS 1.04 0.00 1.03 0.00 1.01 0.00 1.05 0.00 1.08 0.00 1.12 0.00 1.10 0.00 1.07
TVP-QNS 0.79 1.00** 0.79 1.00** 0.79 1.00** 0.84 1.00** 0.86 1.00** 0.87 1.00** 0.79 1.00** 0.83
Notes: Superscripts and ∗∗ indicate inclusion in the model confidence sets $\widehat{\mathcal{M}_{90\%}^{*}}$ and $\widehat{\mathcal{M}_{75\%}^{*}}$ respectively, as per Hansen et al. (2011). The reported average score is computed across all 17 maturities. A dash (’–’) indicates that the quantile scores for the TVP-AR(1)-SV model are exceptionally large and thus omitted from the table for clarity.
Tables 2 and 3 present the 10th and 90th percentile quantile scores (QS) from the out-of-sample forecasting exercise for selected maturities. We also report the corresponding p-values from the Model Confidence Set (MCS) procedure proposed by Hansen et al. (2011). The results suggest that the proposed quantile Nelson-Siegel model delivers superior tail risk forecasts, particularly at longer forecast horizons and for long-term maturities. At shorter horizons, however, the BVAR(1)-CSV model remains competitive and, in some cases, outperforms the proposed model in forecasting tail risks. Importantly, the quantile Nelson-Siegel model consistently outperforms its conditional mean counterparts, based on Diebold et al. (2006), across all forecast horizons and quantile levels. Moreover, the MCS analysis shows that the proposed model is always included in the $\widehat{\mathcal{M}_{90\%}^{*}}$ and $\widehat{\mathcal{M}_{75\%}^{*}}$ confidence sets, especially for longer maturities and extended forecast horizons. By contrast, the three conditional mean models are frequently excluded from these sets. Overall, these findings provide strong empirical support for the proposed time-varying parameter quantile Nelson-Siegel model as a robust tool for forecasting tail risk.
Table 3: Average 90% (Upper-Tail) Quantile Scores for Selected Maturities Relative to an AR(1) Benchmark
Model 3m $p_{\text{MCS}}$ 6m $p_{\text{MCS}}$ 12m $p_{\text{MCS}}$ 24m $p_{\text{MCS}}$ 36m $p_{\text{MCS}}$ 60m $p_{\text{MCS}}$ 120m $p_{\text{MCS}}$ Avg.
1 Month-Ahead Forecast
TVP-AR(1)-SV 0.37 0.00 0.39 0.01 0.49 0.03 0.65 0.00 0.75 0.00 0.85 0.00 0.90 0.00 0.69
BVAR(1)-CSV 0.21 1.00** 0.24 1.00** 0.31 1.00** 0.40 1.00** 0.46 1.00** 0.54 1.00** 0.62 1.00** 0.44
DNS 1.09 0.00 1.04 0.00 1.15 0.00 1.16 0.00 1.12 0.00 1.09 0.00 1.55 0.00 1.17
TVP-QNS 0.53 0.00 0.57 0.00 0.63 0.04 0.59 0.00 0.56 0.02 0.56 0.70** 0.83 0.00 0.61
12 Month-Ahead Forecast
TVP-AR(1)-SV 0.89 1.00** 0.85 1.00** 0.90 1.00** 0.92 1.00** 0.91 1.00** 0.90 1.00** 0.87 1.00** 0.89
BVAR(1)-CSV 0.83 0.93** 0.87 0.93** 0.93 0.93** 0.97 0.93** 0.99 0.94** 1.03 0.93** 1.09 0.94** 0.98
DNS 1.07 0.00 1.08 0.00 1.10 0.00 1.07 0.00 1.03 0.00 0.98 0.00 1.06 0.00 1.04
TVP-QNS 0.99 0.93** 0.99 0.93** 0.98 0.93** 0.90 0.93** 0.83 0.94** 0.76 0.93** 0.77 0.94** 0.86
36 Month-Ahead Forecast
TVP-AR(1)-SV 0.00 0.00 0.00 0.00 0.00 0.00 1.37 0.00
BVAR(1)-CSV 1.68 0.00 1.68 0.00 1.70 0.00 1.72 0.00 1.71 0.00 1.76 0.00 1.90 0.00 1.75
DNS 1.07 0.05 1.06 0.26∗∗ 1.06 0.74∗∗ 1.03 0.05 0.99 0.00 0.96 0.00 1.03 0.00 1.02
TVP-QNS 1.06 1.00** 1.03 1.00** 1.01 1.00** 0.95 1.00** 0.90 1.00** 0.86 1.00** 0.87 1.00** 0.93
Notes: Superscripts and ∗∗ indicate inclusion in the model confidence sets $\widehat{\mathcal{M}_{90\%}^{*}}$ and $\widehat{\mathcal{M}_{75\%}^{*}}$ respectively, as per Hansen et al. (2011). The reported average score is computed across all 17 maturities. A dash (’–’) indicates that the quantile scores for the TVP-AR(1)-SV model are exceptionally large and thus omitted from the table for clarity.
@@ 4.3 In-Sample Results
Previous studies, including Diebold and Li (2006); Diebold et al. (2006, 2008); Mumtaz and Surico (2009); Mönch (2012), focused on the yield curve and its effects within a conditional mean framework. Our study is distinct from the former as it is the first to explore the yield curve from a quantile or distributional perspective.
This section presents our empirical in-sample results. The first subsection reports posterior estimates of the yield curve factors and the decay parameter across various quantile levels. The second subsection investigates the distributional properties of short- and long-term interest rates, as well as selected Treasury maturities, during recent crises—the Great Recession and the COVID-19 pandemic. Finally, the third subsection analyses the relationship between yield curve factors and macroeconomic variables from a quantile perspective.
4.3.1 Posterior Estimates of the Decay Parameter
Figure 2: Posterior distribution of $\lambda(\tau)$ for the 10%, 25%, 50%, 75% and 90% quantiles, with the posterior mean (dashed line).
Figure 2 presents the posterior distribution of the decay parameter $\lambda(\tau)$ across five selected quantiles (10%, 25%, 50%, 75%, and 90%). The results indicate that, for each quantile, the posterior mass of $\lambda(\tau)$ is concentrated in the range of approximately 0.06 to 0.07. This is broadly consistent with the fixed value of $\lambda=0.0609$ proposed by Diebold et al. (2006).
@@ 4.3.2 Quantile Yield Curve Factor Estimates
Figures 3, 4, and 5 depict the posterior mean estimates of time-varying yield curve factors at five quantile levels, providing insights into their dynamic evolution. A consistent finding across all graphs is the presence of substantial temporal variation. Beginning with the level factor, interpreted as the long-term interest rate (i.e., yield as maturity approaches infinity), a steady decline in this rate is discernible across all quantiles since the oil price crisis of the 1980s. Figure 3 reveals a distinct pattern: following both the dot-com recession in the early 2000s and the Great Recession of 2008-09, the posterior estimates for the 90th and median quantiles closely align, while those for the 10th quantile diverge. Conversely, after the COVID-19 recession, the opposite trend emerged, with the 10th and median quantiles exhibiting proximity while the 90th quantile diverged. This evidence suggests that post the early 2000s recession and the Great Recession, the distribution of the long-term interest rate skewed leftwards. In contrast, after the COVID-19 recession, the distribution skewed rightwards, likely attributed to the pronounced inflationary pressures experienced in the aftermath of the COVID-19 period.
@@
Turning to the slope and curvature of the yield curve in Figures 4 and 5, the degree of uncertainty associated with these factors varies over time. Notably, the curvature factor exhibits higher volatility across all quantiles than its slope counterpart. [^1] We impose the monotonicity constraint specified in equation (18) on the short-term interest rate, defined as the sum of the level and slope factors. Figure 10 in the appendix displays the posterior estimates of the short-term rate across the five quantiles, which exhibit no crossing and are thus consistent with the imposed constraint. However, this restriction does not imply monotonicity in the slope factor alone, which may still exhibit crossings across quantiles. In contrast, crossings in the conditional quantiles of the curvature factor are not of concern, as the curvature does not correspond directly to any specific yield maturity. A distinct pattern is evident before 1985, where the posterior estimates for the slope factor across all five quantiles are closely clustered, indicating a period of low uncertainty. In contrast, during the oil price crisis of the 1980s, the curvature factor displayed substantial dispersion across quantiles.
@@
4.3.3 Distributional Characteristics of Treasury Yields Across the Term Structure
The proposed novel framework offers a distinct advantage by enabling the examination of the distributional properties of both short- and long-term interest rates, as well as selected yield maturities, over time. As previously discussed, the long-term interest rate is identified with the level factor, whereas the short-term interest rate reflects a combination of the level and slope factors. We follow the methodology outlined by Mitchell et al. (2024) to construct the empirical distributions of these rates and maturities. Specifically, we adopt an equal-weighting scheme to aggregate all MCMC draws across five quantiles and apply the Epanechnikov kernel, as proposed by Gaglianone and Lima (2012).
@@ 4.3.3 Distributional Characteristics of Treasury Yields Across the Term Structure
Focusing first on the short-term interest rate, we observe clear evidence of a pronounced distributional shift across the three selected months during both the Great Recession and the COVID-19 recession. In particular, following the COVID-19 recession (December 2020), the distribution of the short-term rate becomes less symmetric and exhibits increased negative skewness.
Turning to the long-term interest rate, distributional changes appear more pronounced during the COVID-19 recession than during the Great Recession. After the Great Recession (December 2009), the distribution becomes more negatively skewed. In contrast, during the COVID-19 recession, the distribution shifts towards positive skewness, indicating a different form of market adjustment at the long end of the yield curve.
Figure 7 presents the estimated conditional distributions of three selected Treasury maturities—3 months, 3 years, and 10 years—representing the short, intermediate, and long segments of the yield curve, respectively, across the two crisis periods. Prior to the Great Recession, the distributions of these maturities are relatively similar. However, in the aftermath of the Great Recession, we observe a clear location shift across all maturities, with the 10-year maturity displaying increased left-skewness. During the COVID-19 recession, only the long end of the yield curve undergoes a location shift, accompanied by greater dispersion. Following the COVID-19 episode, all three maturities exhibit a location shift, with the short and intermediate maturities also displaying increased negative skewness.
These results indicate that short- and long-term interest rates, as well as selected yield maturities, experience heightened skewness and asymmetry during recessionary periods. This highlights the empirical relevance of quantile-based approaches, which allow for a more flexible representation of the conditional distribution—particularly in the tails—than traditional factor models such as Nelson-Siegel, which impose symmetry and are limited in their ability to capture such nonlinearities.
Figure 7: Distribution of the 3-month, 3-year, and 10-year Treasury yields during the Great Recession (top panel) and the COVID-19 pandemic period (bottom panel). The solid blue line corresponds to the 3-month maturity, the red dashed line represents the 3-year maturity, and the black dashed line denotes the 10-year maturity.
4.3.4 The Interaction between the Yield Curve and Macroeconomy
The proposed framework, as specified in Equation (9), enables an in-depth analysis of the dynamic interaction between yield curve factors and macroeconomic variables by incorporating a VAR-like state equation. To explore these interactions, we compute the generalized forecast error variance decomposition (GFEVD) across five quantiles, evaluating the contribution of shocks to yield curve factors on macroeconomic variables—and vice versa—as illustrated in Figures 8 and 9, respectively.
We begin by examining the effects of yield curve shocks on macroeconomic variables (Figure 8). The analysis reveals that curvature shocks generate heterogeneous impacts across the distribution of macroeconomic outcomes. Notably, these shocks have a more pronounced effect on the upper tail of all three macroeconomic variables—particularly at longer forecast horizons—indicating heightened sensitivity in more expansionary states of the economy. In contrast, level shocks increasingly influence the lower tail of industrial production over time, suggesting growing downside risks in response to shifts in the overall level of the yield curve.
We next assess the influence of macroeconomic shocks on the yield curve (Figure 9). Shocks to industrial production significantly affect the lower tails of all three yield curve factors, implying that negative real activity shocks disproportionately compress the yield distribution. By contrast, inflation shocks primarily impact the upper tails of the slope and curvature factors, with these effects intensifying at longer horizons—consistent with rising term premia during inflationary episodes. Additionally, federal funds rate shocks increasingly affect the upper tail of the level factor over time, underscoring the asymmetric transmission of monetary policy to the term structure.
Figure 8: Posterior estimates of the generalized forecast error variance decomposition for the macroeconomic variables (columns) on Yield Curve Factors (rows). The colored bars represent different quantiles: blue (10%), orange (25%), yellow (50%), purple (75%), and green (90%).
Figure 9: Posterior estimates of the generalized forecast error variance decomposition for the Yield Curve Factors (columns) on Yield Curve Factors (rows). The colored bars represent different quantiles: blue (10%), orange (25%), yellow (50%), purple (75%), and green (90%).
In summary, these results provide robust evidence that the interaction between yield curve factors and macroeconomic variables is highly dependent on the quantile considered. This highlights the importance of employing quantile-based approaches, which can capture distributional asymmetries and tail behavior that standard linear models may overlook.